OpenStudy (anonymous):

Simplify seventeen s plus three t minus nine minus five t plus nine minus twelve s.

OpenStudy (anonymous):

Well first make that into a number equation. 17s + 3t - 5t + 9 - 12s So you are going to want to combine the like terms. The like terms would be 17s and 12s. So take those and make it 17s + 12s = 29s. The other like terms are 3t and 5t. So take those and do as it says: 3t - 5t = -2t. So in the end that is the farthest you could go because there are no more like terms. So it would be 29s - (-2t) + 9.

OpenStudy (anonymous):

fourteen s minus two t minus nine three s t five s minus two t fourteen s minus t

OpenStudy (anonymous):

my only answers lol

OpenStudy (anonymous):

Oh those are your only answers? Ok let me try that again then. :)

OpenStudy (lgbasallote):

@locogatos2 17 - 12 is not 29

OpenStudy (lgbasallote):

+3t - 5t isn't -(-2t) either

OpenStudy (anonymous):

I wasn't doing 17 - 12, I am now. But before I did 17 + 12.

OpenStudy (anonymous):

And yes I am pretty positive that 3 - 5 = -2. If it doesn't equal that, then what does it equal.

OpenStudy (anonymous):

Lol this is my bros 7th grade homework.

OpenStudy (anonymous):

Ok sorry about that. The answer is 5s - (-2). I missed one of the nines before, and messed up the beginning which led to it being wrong. But I did it right this time. This is how I did it. 17s - 12s = 5s. 3t - 5t = -2t. Then 9 - 9 = 0. So then you have 5s - (-2t).